OpenStudy (christos):

Find the value of the function (f(x+h) - f(x))/ h , h is different than 0 , if f(x)=sqrt(x) Idk How to go about starting this one, I am pretty sure If h is different than 0 then something has to be done about it, but still I dont know

hartnn (hartnn):

can you find f(x+h) first ?

hartnn (hartnn):

to get f(x+h), just replace, 'x' in f(x) by 'x+h'

OpenStudy (christos):

yes when I replace x in f(x) its = f(x+h)

hartnn (hartnn):

so, what you get function f(x+h) as ?

hartnn (hartnn):

replace 'x' in sqrt(x) by 'x+h'

OpenStudy (christos):

sqrt(x+h)

hartnn (hartnn):

yes, now lets take numerator, f(x+h) -f(x) =... ?

OpenStudy (christos):

maybe f(x+h-x) ?

hartnn (hartnn):

no, we have f(x) =sqrt(x) f(x+h)=sqrt(x+h) plug these in f(x+h)-f(x)

OpenStudy (christos):

ok so f(x+h) - sqrt(x)

hartnn (hartnn):

why didn't u plug in for f(x+h) ? :P

OpenStudy (christos):

f(sqrt(x)+h) - sqrt(x) ??

OpenStudy (christos):

sqrt(x) + f(h) - sqrt(x) ?

OpenStudy (christos):

is this good?

hartnn (hartnn):

 got that ^ ?